Fill in the missing fractions.
▯ - 3 / 6 = 3 / 6 ????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given▯- 3 / 6 = 3 / 6
▯=3 / 6 + 3 / 6
▯= (3 + 3) / 6
▯=6 / 6
▯=1

Thanks